If two angles are complementary and one angle is 3 less than twice the other angle, find the measure of the two angles. Write an equation and solve it to find the angles. You must use algebra and show your work.
Answer by arunpaul(104)

let the two angle are = A and B
so Angle A + Angle B = 90 degree
or A+ B = 90
again as per question let A = B + 3
putting the value of A in the above equation we have
or B + 3 + B = 90
or 2B + 3 = 90
or 2B = 90- 3= 87
B = 87/2 = 43.5 degree
so A = 90 - 43.5 = 46.5 degree (Ans)